# Perfect Divisibility and Coloring of Some Bull-Free Graphs

### v3

@@ 2 Notations and Preliminary Results
((1.4 of [8] )) Every bull-free graph can be obtained via substitution from basic bull-free graphs.
Lemma 5.2
((4.3 of [8] )) Let $G$ be a basic bull-free graph. Then for every vertex $v\in V(G)$, either $G[N(v)]$ is perfect or $G[M(v)]$ is perfect.
Lemma 5.3
((6.1 of [2] )) Let $\cal C$} be a class of graphs. If $\cal C$} is hereditary and polynomially $\chi$-bounded with $\chi\leq\omega^{k}$, then ${\cal C}_{s}$ is polynomially $\chi$-bounded with $\chi\leq\omega^{2k+3}$.
@@
Lemma 5.5
Let $G$ be a locally perfect $(P_{6},bull)$-free graph. Then there exists a partition $(X,Y)$ of $V(G)$ such that $G[X]$ and $G[Y]$ are perfect.
Proof. We prove it by induction on $|V(G)|$. It is certain that it holds when $|V(G)|$ is small. Now we proceed with the inductive proof. If $\omega(G)\leq 2$, by Lemma 4.4, then $\chi(G)\leq 4$ and we are done. Therefore, by inductive hypothesis, we may assume that
@@
Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. If $G[M(v)]$ is perfect, then $(N(v),M(v)\cup\{v\})$ is a partition of $V(G)$ which satisfies the Lemma since $G$ is locally perfect. Therefore, we may assume that $G[M(v)]$ is imperfect. By Lemma 3.3, $G[M(v)]$ contains no odd antihole of number of vertices at least 7. So, by Theorem 1.1, $G[M(v)]$ contains a 5-hole as $G$ is $P_{6}$-free. By Lemma 4.2, we have that $G$ has a homogeneous set, say $S$. Recall that $1<|S|<|V(G)|$.
Since $G$ is connected, we have that there exists a vertex which is complete to $S$, and so $G[S]$ is perfect because $G$ is locally perfect. Since $S\neq\mbox{{\rm\O }}$, we may assume that $x\in S$. Let $G^{\prime}=G[V(G)\setminus(S\setminus\{x\})]$. It is easy to verify that $G^{\prime}$ is a locally perfect $(P_{6}$, bull)-free graph and $|V(G^{\prime})|<|V(G)|$. Therefore, by inductive hypothesis, there exists a partition $(X^{\prime},Y^{\prime})$ of $V(G^{\prime})$ such that $G[X^{\prime}]$ and $G[Y^{\prime}]$ is perfect. Since $x\in V(G^{\prime})$, without loss of generality, we may assume that $x\in X^{\prime}$. Let $X=X^{\prime}\cup S$ and $Y=Y^{\prime}$. Clearly, $(X,Y)$ is a partition of $V(G)$. Since $S$ is a homogeneous set of $G$, we have that $S$ is a homogeneous set of $G[X]$, and thus $G[X]$ is a graph obtained from $G[X^{\prime}]$ by substituting the vertex $x$ by the graph $G[S]$. By Lemma 5.4, $G[X]$ is perfect. Therefore, the partition $(X,Y)$ satisfies the lemma. This completes the proof of Lemma 5.5. $\square$
Lemma 5.6
Let $G$ be a basic $(P_{6},bull)$-free graph. Then $\chi(G)\leq\omega(G)^{2}$.
Proof. We prove it by induction on $|V(G)|$. It is certain that it holds when $|V(G)|$ is small. Now we proceed with the inductive proof. We may assume that $\omega(G)\geq 2$ and $G$ is connected by inductive hypothesis.
If $G$ is locally perfectly, then $\chi(G)\leq 2\omega(G)\leq\omega(G)^{2}$ by Lemma 5.5, and so we are done. Therefore, we may assume that $G$ is not locally perfect. Now, by Lemma 5.1, there exists a vertex $v$ such that $G[M(v)]$ is perfect. By inductive hypothesis, we have that $\chi(G[N(v)])\leq(\omega(G)-1)^{2}$. Therefore, we have that
@@
This completes the proof of Lemma 5.6. $\square$
Now, we can prove Theorem 1.5.
Proof of Theorem 1.5: Let $\cal G$} be the class of ($P_{6}$, bull)-free graphs, and $\cal C$} be the class of basic ($P_{6}$, bull)-free graphs. It is easy to verify that $\cal C$} is hereditary. By Lemmas 5.3 and 5.6, it suffices to show that ${\cal G}\subseteq{\cal C}_{s}$.
Let $G\in{\cal G}$. Since $G$ is bull-free, by Lemma 5.1, we have that $G$ can be obtained from a basic bull-free graph $G_{1}$ by substituting a vertex by a basic bull-free graph $G_{2}$. Since $G$ is $P_{6}$-free and $G_{1},G_{2}$ are induced subgraphs of $G$, we have that both $G_{1}$ and $G_{2}$ are $P_{6}$-free, which implies that $G_{1},G_{2}\in{\cal C}$. Therefore, $G\in{\cal C}_{s}$, and so ${\cal G}\subseteq{\cal C}_{s}$.
@@
Remark
[…]
More generally, let ${\cal G}_{1}$ be a hereditary class of graphs in which every trianlge-free graph is $k$-colorable, let ${\cal G}_{2}$ be the class of bull-free graphs, and let ${\cal G}={\cal G}_{1}\cap{\cal G}_{2}$. In [6], Chudnovsky et al. proved that $\cal G$} is polynomially $\chi$-bounded. Here, what are we more concerned about is that is it true that every graph in $\cal G$} is perfectly divisible if and only if $k=3$ ?

### v5

@@ 2 Notations and Preliminary Results
((Lemma 5 of [4])) Let $G$ be a connected (bull, odd net)-free graph, and let $v\in V(G)$. Then $G[M(v)]$ contains no odd antihole of number of vertices at least 7.
@@
We now proceed to prove Theorems 1.5-1.7.
Proof of Theorem 1.5: We prove the Theorem by induction on $|V(G)|$. It is certain that it holds when $|V(G)|$ is small. Now we proceed with the inductive proof. If $\omega(G)\leq 2$, by Lemma 2.5, then $\chi(G)\leq 4$ and we are done. Therefore, by inductive hypothesis, we may assume that
@@
Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. If $G[M(v)]$ is perfect, then $(N(v),M(v)\cup\{v\})$ is a partition of $V(G)$ which satisfies the lemma since $G$ is odd net-free. Therefore, we may assume that $G[M(v)]$ is imperfect. By Lemma 5.4, $G[M(v)]$ contains no odd antihole of number of vertices at least 7. So, by Theorem 1.1, $G[M(v)]$ contains a 5-hole as $G$ is $P_{6}$-free. By Lemma 4.1, we have that $G$ has a homogeneous set, say $S$. Recall that $1<|S|<|V(G)|$.
Since $G$ is connected, we have that there exists a vertex which is complete to $S$, and so $G[S]$ is perfect as $G$ is odd net-free. Since $S\neq\mbox{{\rm\O }}$, we may assume that $x\in S$. Let $G^{\prime}=G[V(G)\setminus(S\setminus\{x\})]$. Therefore, by inductive hypothesis, there exists a partition $(X^{\prime},Y^{\prime})$ of $V(G^{\prime})$ such that $G[X^{\prime}]$ and $G[Y^{\prime}]$ is perfect. Since $x\in V(G^{\prime})$, without loss of generality, we may assume that $x\in X^{\prime}$. Let $X=X^{\prime}\cup S$ and $Y=Y^{\prime}$. Clearly, $(X,Y)$ is a partition of $V(G)$. Since $S$ is a homogeneous set of $G$, we have that $S$ is a homogeneous set of $G[X]$, and thus $G[X]$ is a graph obtained from $G[X^{\prime}]$ by substituting the vertex $x$ by the graph $G[S]$. By Lemma 5.5, $G[X]$ is perfect. Therefore, the partition $(X,Y)$ satisfies the lemma. This completes the proof of Theorem 1.5. $\square$
Proof of Theorem 1.6: We prove the Theorem by induction on $|V(G)|$. It is certain that it holds when $|V(G)|$ is small. Now we proceed with the inductive proof. We may assume that $\omega(G)\geq 2$ and $G$ is connected by inductive hypothesis.
If $G$ is odd net-free, then $\chi(G)\leq 2\omega(G)\leq\omega(G)^{2}$ by Theorem 1.5, and so we are done. Therefore, we may assume that $G$ contains an odd net. Now, by Lemma 5.2, there exists a vertex $v$ such that $G[M(v)]$ is perfect. By inductive hypothesis, we have that $\chi(G[N(v)])\leq(\omega(G)-1)^{2}$. Therefore, we have that
@@
Proof of Theorem 1.7: Let $\cal G$} be the class of (bull, $P_{6}$)-free graphs, and $\cal C$} be the class of (bull, $P_{6}$, odd net+)-free graphs. By Lemma 5.3 and Theorem 1.6, it suffices to show that ${\cal G}\subseteq{\cal C}_{s}$.
Let $G\in{\cal G}$. Since $G$ is bull-free, by Lemma 5.1, we have that $G$ can be obtained from a (bull, odd net+)-free graph $G_{1}$ by substituting a vertex by (bull, odd net+)-free -free graph $G_{2}$. Since $G$ is $P_{6}$-free and $G_{1},G_{2}$ are induced subgraphs of $G$, we have that both $G_{1}$ and $G_{2}$ are $P_{6}$-free, which implies that $G_{1},G_{2}\in{\cal C}$. Therefore, $G\in{\cal C}_{s}$, and so ${\cal G}\subseteq{\cal C}_{s}$.
This completes the proof of Theorem 1.7. $\square$